The answer to each of the questions is a single digit integer ranging from 0 to 9.


 






The form of the dependence of the light force per unit wingspan F (wingspan is the total length of the wings) on an aircraft wing of width L, moving with a velocity v through air of density r is F = kxLyvzrw. Find the value of x + y + z + w.



 

Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

6

 

        F is lift force per unit wingspan i.e. lift force per unit length.

 

        So [F] =  = MT–2

 

        F µ Lavbpc Þ F = kLavbpc

 

        Þ MT–2 = La(LT–1)b(ML–3)c

 

        Hence a + b – 3c = 0, c = 1, –b = –2 ® b = 2, a = 1

 

        ® F = k L2v2p.

 

        \ x + y + z + w = 1 + 2 + 2 + 1 = 6